OpenStudy (anonymous):

How many different 9 player teams can be fielded if there are 3 players who can only play catcher, 4 player who can only play first base , 6 player who can only pitch ,and 14 players who can play in any in any of the remain 6 position?

OpenStudy (anonymous):

\( \binom{3}{1} \times \binom{4}{1} \times \binom{6}{1} \times \binom{14}{6} = \cdots \)

OpenStudy (anonymous):

not match the key

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I solve your equation not match with the key is 216,216

OpenStudy (anonymous):

answer is 216,216 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes, that's correct. \[ \binom{3}{1} \times \binom{4}{1} \times \binom{6}{1} \times \binom{14}{6} = 216216 \]

OpenStudy (anonymous):

3C1*4C1*6C1*14C6 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

3* 4* 6* 3003 = 216,216 OK IT correct thank

OpenStudy (anonymous):

glad to help :)